Prove that Z[√2] is isomorphic to

Matrix H = a 2b

b a

Where a,b∈Z as rings.


1
Expert's answer
2021-06-08T04:44:17-0400

"Let \\Z[
√
2] =\{ a + b
√
2|a, b ∈ Z\} and"

"H=\{ \begin{bmatrix} a&2b\\ b&a \end{bmatrix} | a,b \isin Z \}"

show that Z[√2] and Hare isomorphic as rings


solution.

define a function

"f: H\to Z[\sqrt{2}] by \\
f\begin{pmatrix}\begin{bmatrix}a&2b\\b&a\end{bmatrix}\end{pmatrix}=a+b\sqrt{2}\\
then, for A=\begin{bmatrix}a&2b\\b&a\end{bmatrix},
C=\begin{bmatrix}c&2d\\d&c\end{bmatrix}\isin H, we have\\
f\begin{pmatrix}\begin{bmatrix}a&2b\\b&a\end{bmatrix}+\begin{bmatrix}c&2d\\d&c\end{bmatrix}\end{pmatrix}=\\
f\begin{pmatrix}\begin{bmatrix}a+c&2b+2d\\b+d&a+c\end{bmatrix}\end{pmatrix}\\=
(a+b)+(b+d)\sqrt{2}\\
=(a+b\sqrt{2})+(c+d\sqrt{2})\\
=f\begin{pmatrix}\begin{bmatrix}a&2b\\b&a\end{bmatrix}\end{pmatrix}+
f\begin{pmatrix}\begin{bmatrix}c&2d\\d&c\end{bmatrix}\end{pmatrix}\\
and\\
f\begin{pmatrix}\begin{bmatrix}a&2b\\b&a\end{bmatrix}.\begin{bmatrix}c&2d\\d&c\end{bmatrix}\end{pmatrix}=\\
f\begin{pmatrix}\begin{bmatrix}ac+2bd&2ad+2bc\\ad+bc&ac+2bd\end{bmatrix}\end{pmatrix}=(ac+2bd)+(ad+bc)\sqrt{2}\\
=(a+b\sqrt{2})(c+d\sqrt{2})\\
=f\begin{pmatrix}\begin{bmatrix}a&2b\\b&a\end{bmatrix}\end{pmatrix}.f\begin{pmatrix}\begin{bmatrix}c&2d\\d&c\end{bmatrix}\end{pmatrix}
\\"


Hence, f is a ring homomorphism. Now, f is surjective since, given any x ∈ Z[√2], we have that x = a + b√2 for some a, b ∈ Z, so A ∈ H and f(A)=a+b√2=x.

finally to see that f is injective

suppose that A,C ∈ H and f(A)=f(C)

this mean that a+b√2=c+d√2.

if b not equal to d (so that d-b is not equal to zero). then this would imply that √2 =(a-c)/(d-b) is rational, a contradiction. hence we must have that b=d.

hence we have

a + b√2 = c + d√2 = c + b√2.

Subtracting b√2 from both sides of the above equation shows that a = c. Thus,

A=C

, so it follows that f is injective. Hence, f is a ring isomorphism.